OpenStudy (anonymous):

Write the equation y - 2 = 4(x + 5) in standard form. A. 4x + y = 7 B. 4x - y = 3 C. 4x - y = -22

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (nurali):

y - 2 = 4(x + 5) y = 4(x + 5)+2 y = 4x + 20+2 y=4x+22 4x-y=-22